solve the following system of equations by the addition method
5x - y=6
-2x + 3y=5
what is the value of the x-coordinate?

Answer by fractalier(6550)

From
5x - y = 6
-2x + 3y = 5
let us multiply the top equation by three and add it to the second...
-2x + 3y = 5
15x - 3y = 18
------------------
13x = 23
x = 23/13